store sells bluegrass seed that is worth $6 per pound and ryegrass seed that is worth $3 per pound. How much of each should be mixed to obtain 80 pounds of a blend that is worth $5 per pound?
Answer by stanbon(75887)

store sells bluegrass seed that is worth $6 per pound and ryegrass seed that is worth $3 per pound. How much of each should be mixed to obtain 80 pounds of a blend that is worth $5 per pound?
-------------
Weight Equation:: b + r = 80 lbs
Value Equation::: 6b + 3r = 5*80 dollars
----------------
Modify::
6b + 6r = 6*80
6b + 3r = 5*80
----
3r = 80
r = 80/3 = 26 2/3 lbs (amt. of rye seed needed)
-----
Solve for "b"::
b + (80/3) = 80
b = 53 1/3 lb (amt. of bluegrass seed needed)
-----------
Cheers,
Stan H.
--------------